### v1

@@ Abstract
Neural Combinatorial Optimization (NCO) has emerged as a promising approach for NP-hard problems. However, prevailing RL-based methods suffer from low sample efficiency due to sparse rewards and underused solutions. We propose Preference Optimization for Combinatorial Optimization (POCO), a training paradigm that leverages solution preferences via objective values. It introduces: (1) an efficient preference pair construction for better explore and exploit solutions, and (2) a novel loss function that adaptively scales gradients via objective differences, removing reliance on reward models or reference policies. Experiments on Job-Shop Scheduling (JSP), Traveling Salesman (TSP), and Flexible Job-Shop Scheduling (FJSP) show POCO outperforms state-of-the-art neural methods, reducing optimality gaps impressively with efficient inference. POCO is architecture-agnostic, enabling seamless integration with existing NCO models, and establishes preference optimization as a principled framework for combinatorial optimization.
@@ 1 Introduction
Combinatorial optimization problems (COPs), such as scheduling Zhang et al. (2019); Xiong et al. (2022) and routing problems Vidal et al. (2020); Berghman et al. (2023), are widely applied in real-world scenarios and have attracted significant research attention. Most COPs are NP-hard, making them challenging to find optimal solutions. Exact methods, such as branch-and-bound algorithms, require exponential computation time as the problem size increases. Consequently, heuristic methods have proven effective in obtaining high-quality solutions within reasonable time over the past decades. Nevertheless, these methods still heavily rely on expert knowledge and extensive iterative search.
@@
In the emerging field of neural combinatorial optimization (NCO), deep neural models are employed to automatically learn heuristics from training data, enabling the rapid construction of high-quality solutions in an end-to-end fashion Mazyavkina et al. (2021); Bengio et al. (2021); Yan et al. (2022); Kayhan & Yildiz (2023); Zhang et al. (2023); Garmendia et al. (2024). Early research Vinyals et al. (2015) adopted supervised learning (SL) to train the deep models, which required (near-) optimal solutions produced by expensive specialized solvers as labels. Different from SL, reinforcement learning (RL), which does not require labels, has emerged as the mainstream training paradigm for NCO Kool et al. (2019); Kwon et al. (2020); Zhang et al. (2020); Ho et al. (2024). However, RL encounters challenges such as sparse rewards and low sample efficiency Kim et al. (2024). Recently, self-labeling learning (SLL) Corsini et al. (2024); Luo et al. (2024) was proposed to partially address these issues by sampling multiple solutions and treating the best one among them as a pseudo-label for model training. Nevertheless, SLL still faces the challenge of low sample efficiency, as all sampled solutions except the optimal one are discarded during training.
@@
To improve sample efficiency in NCO training, we leverage multiple sampled solutions rather than focusing solely on the optimal one by introducing preference optimization Rafailov et al. (2024); Meng et al. (2024). To this end, we propose Preference Optimization for Combinatorial Optimization (POCO), which leverages the natural preference relation among solutions according to their objective values. Our approach builds upon two fundamental observations: (1) NCO models (typically generative models) can generate multiple distinct solutions for a given problem instance, and (2) the objective value of a COP solution can be computed with a low cost. As shown in Figure 1, our POCO comprises two essential components: constructing multiple preference pairs from sampled solutions and building a preference optimization loss. As a new training paradigm for NCO, POCO avoids expensive labels and formulation of Markov Decision Process, making it particularly well-suited for addressing various types of COPs.
Figure 1: The pipeline of preference optimization for combinatorial optimization (POCO).
@@
  • We propose POCO, a novel training paradigm using preference optimization for NCO, which enhances sample efficiency compared with the mainstream RL and recent SLL paradigms.
  • As the first key component of POCO, we design a preference pair construction method for COPs to better explore and exploit solutions.
  • As the second key component, we tailor a novel preference optimization loss that incorporates preferences quantified by objective values of COPs.
  • Experimental results on three classic COPs, namely the Job-shop Scheduling Problem (JSP), Traveling Salesman Problem (TSP), and Flexible Job-shop Scheduling Problem (FJSP), demonstrate that POCO outperforms state-of-the-art NCO methods.
@@ 2 Related Works
Supervised Learning (SL) for NCO. SL methods utilize optimal solutions as labels to train neural models with cross-entropy loss for solving COPs, such as TSP Vinyals et al. (2015); Milan et al. (2017) and JSP Ingimundardottir & Runarsson (2018). Data augmentation techniques have been exploited to enhance the performance of SL methods for routing problems Luo et al. (2023); Yao et al. (2024). Additionally, diffusion-based SL approaches Sun & Yang (2023); Li et al. (2023; 2024) learn to generate heatmaps for TSP. The primary limitation of SL, however, lies in the high computational cost of obtaining optimal solutions as labels, which restricts its practical applications.
@@
Reinforcement Learning (RL) & Self-Labeling Learning (SLL) for NCO. Label-free RL is currently the mainstream training paradigm in the NCO field. Attention Model Kool et al. (2019), which combines RL with Transformer Vaswani (2017), marks a milestone in solving routing problems. The policy optimization with multiple optima (POMO) Kwon et al. (2020) introduces a shared baseline leveraging solution symmetry. Given its competitive performance and practicality, POMO has established itself as a prominent training algorithm for routing problems and has inspired numerous advancements Grinsztajn et al. (2023); Chalumeau et al. (2023); Drakulic et al. (2023); Chen et al. (2023a; b); Xiao et al. (2024b); Goh et al. (2024); Fang et al. (2024); Zhou et al. (2024a; b); Bi et al. (2024); Zheng et al. (2024); Wang et al. (2024). Additionally, RL has been widely adopted for scheduling problems, including JSP Zhang et al. (2020); Park et al. (2021a; b); Jeon et al. (2023); Tassel et al. (2023); Iklassov et al. (2023); Ho et al. (2024) and FJSP Song et al. (2023); Yuan et al. (2024). Different from the above RL-based constructive methods, RL is also employed in improvement methods for both routing Ma et al. (2021); Wu et al. (2022); Ma et al. (2023); Kong et al. (2024) and scheduling problems Falkner et al. (2022); Zhang et al. (2024b). Recently, SLL Luo et al. (2024); Corsini et al. (2024) utilizes the local optimal solution during training as a pseudo-label to train an end-to-end model using cross-entropy loss, where self-labeling improvement method (SLIM) Corsini et al. (2024) has achieved state-of-the-art performance on JSP.
@@
Preference Optimization. Preference optimization has been widely adopted to align large language models (LLMs) with human preferences. One of the most well-known techniques is reinforcement learning with human feedback (RLHF) Stiennon et al. (2020); Ouyang et al. (2022), which trains a reward model using ranking learning and then aligns LLM through RL. Direct preference optimization (DPO) Rafailov et al. (2024) offers an efficient alternative by skipping the reward model training phase and directly optimizing LLMs using preference pairs. Building upon DPO, subsequent studies explore comparing more samples Dong et al. (2023); Song et al. (2024) and designing more concise loss functions Xu et al. (2023); Meng et al. (2024). Among them, simple preference optimization (SimPO) Meng et al. (2024) has gained popularity due to its simplicity and effectiveness. Recent advancements in preference optimization are summarized in a comprehensive survey Xiao et al. (2024a). Inspired by preference optimization, we propose preference optimization for combinatorial optimization (POCO), in which we develop a preference pair construction method and a novel preference optimization loss specially designed for COPs.
@@ 3.1 Neural Combinatorial Optimization (NCO)
Typical training paradigms include SL, RL, and SLL. SL utilizes the (near-) optimal solution $\bm{y}^{*}$ as a label to train model $\bm{\theta}$ using cross-entropy loss $\mathcal{L}(\pi_{\bm{\theta}}|\bm{y}^{*},\bm{x})=-\log\pi_{\bm{\theta}}(\bm{y}^{*}|\bm{x})$. In RL, the REINFORCE loss Williams (1992) $\mathcal{L}(\pi_{\bm{\theta}}|\bm{y},\bm{x})=-(g(\bm{y})-b(\bm{x}))\log\pi_{\bm{\theta}}(\bm{y}|\bm{x})$ with a baseline $b(\bm{x})$ is commonly used for routing problems, while the proximal policy optimization (PPO) Schulman et al. (2017) loss $\mathcal{L}(\pi_{\bm{\theta}}|\bm{y},\bm{x})=-\sum_{t}\min\{r_{t}\hat{A}_{t},\text{clip}(r_{t},1-\epsilon,1+\epsilon)\hat{A}_{t}\}$ is predominantly used for scheduling problems, where $r_{t}=\frac{\pi_{\bm{\theta}}(y_{t}|\bm{y}_{<t},\bm{x})}{\pi_{{\bm{\theta}}^{\prime}}(y_{t}|\bm{y}_{<t},\bm{x})}$ denotes the ratio of current policy $\bm{\theta}$ and old policy ${\bm{\theta}}^{\prime}$, $\hat{A}_{t}$ represents the advantage estimate, and $\epsilon$ is a hyperparameter. SLL, a recent paradigm, selects the best among solutions sampled from current policy $\bm{\theta}$ as a pseudo-label and applies cross-entropy loss.
@@ 4 Methodology
4.1 Preference Optimization for Combinatorial Optimization (POCO)
To improve sample efficiency, we propose a novel training paradigm, namely POCO. Distinct from the RL and SLL training paradigms, POCO exploits the preference relations among generated solutions according to their objective values. Specifically, for a COP with the minimization objective, the explicit preference $f^{*}(\bm{y},\bm{x})$ for instance $\bm{x}$ and solution $\bm{y}$ is defined as the negative of the objective function:
@@ 4.1 Preference Optimization for Combinatorial Optimization (POCO)
POCO employs a preference optimization loss based on such preference pairs to train the NCO model parameterized as $\bm{\theta}$. As the two critical components of POCO, we develop a preference pair construction method and derive a novel loss function specialized for COPs.
4.2 Preference Pair Construction
@@
4.3 Loss Function
After obtaining preference pairs, we formulate the loss function of POCO by incorporating a preference-based scaling factor derived from the objective values of COP solutions.
POCO Loss. For policy $\pi_{\bm{\theta}}(\bm{y},\bm{x})$ used to construct solution $\bm{y}$, its implicit preference $f_{\bm{\theta}}(\bm{y},\bm{x})$ is defined as the average log-likelihood:
@@ 4.3 Loss Function
For a preference pair $(\bm{x},\bm{y}_{w},\bm{y}_{l})$, the preference distribution $p_{\bm{\theta}}(\bm{y}_{w}\succ\bm{y}_{l}|\bm{x})$ is modeled using the Bradley-Terry ranking objective Bradley & Terry (1952) and implicit preferences:
@@
By maximizing the log-likelihood of $p_{\bm{\theta}}(\bm{y}_{w}\succ\bm{y}_{l}|\bm{x})$, the model is encouraged to assign higher probabilities to preferred solutions $\bm{y}_{w}$ compared with less preferred solutions $\bm{y}_{l}$. From Equations 1, 2 and 3, we can derive the POCO loss function:
$\mathcal{L}_{POCO}(\pi_{\bm{\theta}},\bm{x},\bm{y}_{w},\bm{y}_{l})=$ (4)
@@
The POCO training algorithm is presented in Algorithm 1.
Algorithm 1 POCO Training
@@
Compute $\mathcal{L}_{POCO}(\pi_{\bm{\theta}},\bm{x},\bm{y}_{w},\bm{y}_{l})$ using Equation 4
@@
$\mathcal{L}(\bm{\theta})\leftarrow\frac{1}{D}\sum_{i=1}^{D}\frac{1}{|\mathcal{P}_{i}|}\sum_{(\bm{x},\bm{y}_{w},\bm{y}_{l})\in\mathcal{P}_{i}}\mathcal{L}_{POCO}(\pi_{\bm{\theta}},\bm{x},\bm{y}_{w},\bm{y}_{l})$
@@
Comparison with Other Losses. Our POCO loss differs from existing preference optimization losses in several key aspects. Compared with RLHF Stiennon et al. (2020), it eliminates the need to train an additional reward model. Compared with DPO Rafailov et al. (2024), it avoids using a reference model, reducing computational costs. Compared with SimPO Meng et al. (2024), it incorporates a preference-based scaling factor without requiring extra hyperparameters, avoiding labor-intensive hyperparameter tuning. Detailed analyses are provided in Appendix C.
4.4 Characteristics of POCO
In summary, our POCO has the following characteristics. (1) Novel training paradigm: POCO introduces preference optimization to NCO as a new training paradigm, featuring two effective problem-awareness components: preference pair construction and POCO loss. (2) Architecture agnostic: POCO is compatible with various NCO models for different COPs, achieving high sample efficiency without expensive labels while inheriting the fast inference advantage of NCO models. (3) Outstanding performance: Compared with existing NCO methods, POCO surpasses state-of-the-art results on classic COPs, including JSP, TSP, and FJSP.

### v3

@@ Abstract
Neural Combinatorial Optimization (NCO) has emerged as a promising approach for NP-hard problems. However, prevailing RL-based methods suffer from low sample efficiency due to sparse rewards and underused solutions. We propose Best-anchored and Objective-guided Preference Optimization (BOPO), a training paradigm that leverages solution preferences via objective values. It introduces: (1) a best-anchored preference pair construction for better explore and exploit solutions, and (2) an objective-guided pairwise loss function that adaptively scales gradients via objective differences, removing reliance on reward models or reference policies. Experiments on Job-shop Scheduling Problem (JSP), Traveling Salesman Problem (TSP), and Flexible Job-shop Scheduling Problem (FJSP) show BOPO outperforms state-of-the-art neural methods, reducing optimality gaps impressively with efficient inference. BOPO is architecture-agnostic, enabling seamless integration with existing NCO models, and establishes preference optimization as a principled framework for combinatorial optimization.
Neural Combinatorial Optimization, Preference Optimization, Machine Learning
@@ 1 Introduction
Combinatorial optimization problems (COPs), such as scheduling (Zhang et al., 2019; Xiong et al., 2022) and routing problems (Vidal et al., 2020; Berghman et al., 2023), are widely applied in real-world scenarios and have attracted significant research attention. Most COPs are NP-hard, making them challenging to find optimal solutions. Exact methods, such as branch-and-bound algorithms, require exponential computation time as the problem size increases. Consequently, heuristic methods have proven effective in obtaining high-quality solutions within reasonable time over the past decades. Nevertheless, these methods still heavily rely on expert knowledge and extensive iterative search.
@@
In the emerging field of neural combinatorial optimization (NCO), deep neural models are employed to automatically learn heuristics from training data, enabling the rapid construction of high-quality solutions in an end-to-end fashion (Mazyavkina et al., 2021; Bengio et al., 2021; Yan et al., 2022; Kayhan & Yildiz, 2023; Zhang et al., 2023; Garmendia et al., 2024). Early research (Vinyals et al., 2015) adopted supervised learning (SL) to train the deep models, which required (near-) optimal solutions produced by expensive specialized solvers as labels. Different from SL, reinforcement learning (RL), which does not require labels, has emerged as the mainstream training paradigm for NCO (Kool et al., 2019; Kwon et al., 2020; Zhang et al., 2020; Ho et al., 2024). However, RL encounters challenges such as sparse rewards and low sample efficiency (Kim et al., 2024). Recently, self-labeling learning (SLL) (Corsini et al., 2024; Luo et al., 2025) was proposed to partially address these issues by sampling multiple solutions and treating the best one among them as a pseudo-label for model training. Nevertheless, SLL still faces the challenge of low sample efficiency, as all sampled solutions except the optimal one are discarded during training.
@@
To improve sample efficiency in NCO training, we leverage multiple sampled solutions rather than focusing solely on the optimal one by introducing preference optimization (Rafailov et al., 2023; Meng et al., 2024). To this end, we propose Best-anchored and Objective-guided Preference Optimization (BOPO), which leverages the natural preference relation among solutions according to their objective values. Our approach builds upon two fundamental observations: (1) NCO models (typically generative models) can generate multiple distinct solutions for a given problem instance, and (2) the objective value of a COP solution can be computed with a low cost. As shown in Figure 1, our BOPO comprises two essential components: constructing multiple preference pairs from sampled solutions and building a preference optimization loss. As a new training paradigm for neural combinatorial optimization, BOPO avoids expensive labels and formulation of Markov Decision Process, making it particularly well-suited for addressing various types of COPs.
Figure 1: The pipeline of best-anchored and objective-guided preference optimization (BOPO).
@@
  • We propose BOPO, a novel training paradigm using preference optimization for neural combinatorial optimization, which enhances sample efficiency compared with the mainstream RL and recent SLL paradigms.
  • As the first key component of BOPO, we design a best-anchored preference pair construction method for COPs to better explore and exploit solutions.
  • As the second key component, we tailor a novel objective-guided preference optimization loss that incorporates preferences quantified by objective values of COPs.
  • Experimental results on three classic problems, namely the Job-shop Scheduling Problem (JSP), Traveling Salesman Problem (TSP), and Flexible Job-shop Scheduling Problem (FJSP), demonstrate that BOPO outperforms state-of-the-art methods.
@@ 2 Related Works
Supervised Learning (SL) for NCO. SL methods utilize optimal solutions as labels to train neural models with cross-entropy loss for solving COPs, such as TSP (Vinyals et al., 2015; Milan et al., 2017) and JSP (Ingimundardottir & Runarsson, 2018). Data augmentation techniques have been exploited to enhance the performance of SL methods for routing problems (Luo et al., 2023; Yao et al., 2024). Additionally, diffusion-based SL approaches (Sun & Yang, 2023; Li et al., 2023, 2024) learn to generate heatmaps for TSP. The primary limitation of SL, however, lies in the high computational cost of obtaining optimal solutions as labels, which restricts its practical applications.
@@
Reinforcement Learning (RL) & Self-Labeling Learning (SLL) for NCO. Label-free RL is currently the mainstream training paradigm in neural combinatorial optimization. Attention Model (Kool et al., 2019), which combines RL with Transformer (Vaswani, 2017), marks a milestone in solving routing problems. The policy optimization with multiple optima (POMO) (Kwon et al., 2020) introduces a shared baseline leveraging solution symmetry. Given its competitive performance and practicality, POMO has established itself as a prominent training algorithm for routing problems and has inspired numerous advancements (Grinsztajn et al., 2023; Chalumeau et al., 2023; Drakulic et al., 2023; Chen et al., 2023a, b; Xiao et al., 2024b; Goh et al., 2024; Fang et al., 2024; Zhou et al., 2024a, b; Bi et al., 2024; Zheng et al., 2024; Wang et al., 2024; Chen et al., 2025a, b). Additionally, RL has been widely adopted for scheduling problems, including JSP (Zhang et al., 2020; Park et al., 2021a, b; Jeon et al., 2023; Tassel et al., 2023; Iklassov et al., 2023; Ho et al., 2024) and FJSP (Song et al., 2023; Yuan et al., 2024). Different from the above RL-based constructive methods, RL is also employed in improvement methods for both routing (Ma et al., 2021; Wu et al., 2022; Ma et al., 2023; Kong et al., 2024) and scheduling problems (Falkner et al., 2022; Zhang et al., 2024b). Recently, SLL (Luo et al., 2025; Corsini et al., 2024) utilizes the local optimal solution during training as a pseudo-label to train an end-to-end model using cross-entropy loss, where self-labeling improvement method (SLIM) (Corsini et al., 2024) has achieved state-of-the-art performance on JSP.
@@
Preference Optimization. Preference optimization has been widely adopted to align large language models (LLMs) with human preferences. One of the most well-known techniques is reinforcement learning with human feedback (RLHF) (Stiennon et al., 2020; Ouyang et al., 2022), which trains a reward model using ranking learning and then aligns LLM through RL. Direct preference optimization (DPO) (Rafailov et al., 2023) offers an efficient alternative by skipping the reward model training phase and directly optimizing LLMs using preference pairs. Building upon DPO, subsequent studies explore comparing more samples (Dong et al., 2023; Song et al., 2024) and designing more concise loss functions (Xu et al., 2023; Meng et al., 2024). Among them, simple preference optimization (SimPO) (Meng et al., 2024) has gained popularity due to its simplicity and effectiveness. Recent advancements in preference optimization are summarized in a comprehensive survey (Xiao et al., 2024a). Inspired by preference optimization, we propose best-anchored and objective-guided preference optimization (BOPO). A concurrent work (Pan et al., 2025) also applies preference optimization to COPs. However, fundamentally different from Pan et al. (2025), BOPO develops a best-anchored preference pair construction method and a novel objective-guided preference optimization loss specially designed for COPs.
@@ 3.1 Neural Combinatorial Optimization (NCO)
Typical training paradigms include SL, RL, and SLL. SL utilizes the (near-) optimal solution $\bm{y}^{*}$ as a label to train model $\bm{\theta}$ using cross-entropy loss $\mathcal{L}(\pi_{\bm{\theta}}|\bm{y}^{*},\bm{x})=-\log\pi_{\bm{\theta}}(\bm{y}^{*}|\bm{x})$. In RL, the REINFORCE loss (Williams, 1992) $\mathcal{L}(\pi_{\bm{\theta}}|\bm{y},\bm{x})=-(g(\bm{y})-b(\bm{x}))\log\pi_{\bm{\theta}}(\bm{y}|\bm{x})$ with a baseline $b(\bm{x})$ is commonly used for routing problems, while the proximal policy optimization (PPO) (Schulman et al., 2017) loss $\mathcal{L}(\pi_{\bm{\theta}}|\bm{y},\bm{x})=-\sum_{t}\min\{r_{t}\hat{A}_{t},\text{clip}(r_{t},1-\epsilon,1+\epsilon)\hat{A}_{t}\}$ is predominantly used for scheduling problems, where $r_{t}=\frac{\pi_{\bm{\theta}}(y_{t}|\bm{y}_{<t},\bm{x})}{\pi_{{\bm{\theta}}^{\prime}}(y_{t}|\bm{y}_{<t},\bm{x})}$ denotes the ratio of current policy $\bm{\theta}$ and old policy ${\bm{\theta}}^{\prime}$, $\hat{A}_{t}$ represents the advantage estimate, and $\epsilon$ is a hyperparameter. SLL, a recent paradigm, selects the best among solutions sampled from current policy $\bm{\theta}$ as a pseudo-label and applies cross-entropy loss.
@@ 4 Methodology
4.1 Best-anchored and Objective-guided Preference Optimization (BOPO)
To improve sample efficiency, we propose a novel training paradigm, namely BOPO. Distinct from the RL and SLL training paradigms, BOPO exploits the preference relations among generated solutions according to their objective values. Specifically, for a COP with the minimization objective, the explicit preference $f^{*}(\bm{y},\bm{x})$ for instance $\bm{x}$ and solution $\bm{y}$ is defined as the negative of the objective function:
@@ 4.1 Best-anchored and Objective-guided Preference Optimization (BOPO)
BOPO employs a preference optimization loss based on such preference pairs to train the neural model parameterized as $\bm{\theta}$. As the two critical components of BOPO, we develop a best-anchored preference pair construction method and derive a novel objective-guided preference optimization loss function specialized for COPs.
4.2 Best-anchored Preference Pair Construction
@@
4.3 Objective-guided Preference Optimization Loss
After obtaining preference pairs, we formulate the loss function of BOPO by incorporating a preference-based scaling factor derived from the objective values of COP solutions.
Objective-guided Preference Optimization Loss. For policy $\pi_{\bm{\theta}}(\bm{y},\bm{x})$ used to construct solution $\bm{y}$, its implicit preference $f_{\bm{\theta}}(\bm{y},\bm{x})$ is defined as the average log-likelihood:
@@ 4.3 Objective-guided Preference Optimization Loss
For a preference pair $(\bm{x},\bm{y}_{w},\bm{y}_{l})$, the preference distribution $p_{\bm{\theta}}(\bm{y}_{w}\succ\bm{y}_{l}|\bm{x})$ is modeled using the Bradley-Terry ranking objective (Bradley & Terry, 1952) and implicit preferences:
@@
By maximizing the log-likelihood of $p_{\bm{\theta}}(\bm{y}_{w}\succ\bm{y}_{l}|\bm{x})$, the model is encouraged to assign higher probabilities to preferred solutions $\bm{y}_{w}$ compared with less preferred solutions $\bm{y}_{l}$. From Equations 1, 2 and 3, we can derive the BOPO loss function:
$\mathcal{L}_{BOPO}(\pi_{\bm{\theta}},\bm{x},\bm{y}_{w},\bm{y}_{l})=$ (4)
@@
The BOPO training algorithm is presented in Algorithm 1.
Algorithm 1 BOPO Training
@@
Compute $\mathcal{L}_{BOPO}(\pi_{\bm{\theta}},\bm{x},\bm{y}_{w},\bm{y}_{l})$ using Equation 4
@@
$\mathcal{L}(\bm{\theta})\leftarrow\frac{1}{D}\sum_{i=1}^{D}\frac{1}{|\mathcal{P}_{i}|}\sum_{(\bm{x},\bm{y}_{w},\bm{y}_{l})\in\mathcal{P}_{i}}{\mathcal{L}_{BOPO}}(\pi_{\bm{\theta}},\bm{x},\bm{y}_{w},\bm{y}_{l})$
@@
Comparison with Other Losses. Our BOPO loss differs from existing preference optimization losses in several key aspects. Compared with RLHF (Stiennon et al., 2020), it eliminates the need to train an additional reward model. Compared with DPO (Rafailov et al., 2023), it avoids using a reference model, reducing computational costs. Compared with SimPO (Meng et al., 2024), it incorporates a objective-guided scaling factor without requiring extra hyperparameters, avoiding labor-intensive hyperparameter tuning. Detailed analyses are provided in Appendix C.
Comparison with RL. By leveraging preference learning between the best solution and diverse inferior ones, BOPO guides the model toward promising decision trajectories and discern suboptimal choices, demonstrating its advantage against RL methods. Meanwhile, recognizing that precise rewards (i.e., objective values) in RL are crucial to combinatorial optimization, BOPO introduces an objective-guided scaling factor beyond the standard preference optimization loss function to better distinguish preference differences.
4.4 Characteristics of BOPO
In summary, our BOPO has the following characteristics. (1) Novel training paradigm: BOPO introduces preference optimization to neural combinatorial optimization as a new training paradigm, featuring two effective problem-awareness components: best-anchored preference pair construction and objective-guided preference optimization loss. (2) Architecture agnostic: BOPO is compatible with various models for different problems, achieving high sample efficiency without expensive labels while inheriting the fast inference advantage of neural models. (3) Outstanding performance: Compared with existing methods, BOPO surpasses state-of-the-art results on classic COPs, including JSP, TSP, and FJSP.